Lowered tower from about 11:00-12:00 to replace scrubber bottles in all EC150s.  h2o and co2 values should change a lot over the next 24 hours, hopefully settling on something reasonable tomorrow.  We didn't do a zero/span calibration, as recommended by Campbell as part of this procedure.

Also checked Tsoil.  We could see the "handle" exposed, but not the 0.6cm probe.  We added a bit of dirt to cover both, just in case. 

Nothing obviously concerning about the TP01.

Note that there are lots of wasps in some of the tower structure.  We need to plug a bunch of holes – at least 10 that are a bit less than 1/2" diameter.

We didn't add the "stop" that Dan suggested to prevent the tower height from sagging.